Ben is 4
times as old as Ishaan and is also 6 years older than Ishaan.

Answer by ikleyn(52787)

.
Let x be the Ishaan's age.


Then Ben's age is 4x, and from the problem's description,


    4x = x + 6.


It implies


    4x - x = 6,

    3x     = 6

     x     = 6/3 = 2.


ANSWER.  Ishaan is 2 years old;  Ben is 2*4 = 8 years old.